OpenStudy (anonymous):

Please help! if θ is an acute angle and sin θ= 1/2, then cos θ= A. -1 B. 0 C. 1/2 D. √3/2 E. 2

OpenStudy (xmissalycatx):

You'd basically find the inverse of it so I THINK it is E, but I could be wrong.

OpenStudy (anonymous):

Okay I think So too... thanks :)

jhonyy9 (jhonyy9):

yes @xMissAlyCatx so i think it same wrong your answer because where sinß = 1/2 so this mean that ß = 30° so what will resulted for cosß = sqrt3 /2 - so i think that from this result D. answer will be correct hope this is understandably and will help you please correct your answer @Kfsv10 hope this will help you understanding

OpenStudy (xmissalycatx):

Well like I said @jhonyy9 , I wasn't sure... soo..